Two equal metal balls are charged to 10 and -20 units of electricity. Then they are brought in contact with


each other and then again separated to the original distance. The ratio of magnitudes of the forces between


the two balls is:


(a) + 8 : 1 (b) 1 : 8 (c) +2 : 1 (d) 1 : 2


Grade:11

1 Answers

Anirban Mukherjee
11 years ago

Initial Force =10X20/x^2

Final force=5X5/x^2

Charge=[10-20]/2=-5

F1/F2=200/25=8